# Parallel Heuristic Exploration for Additive Complexity Reduction in Fast Matrix Multiplication

### v1

@@ Abstract
This paper presents a parallel random-search method for reducing additive complexity in fast matrix multiplication. The approach replaces expensive exact evaluation with fast heuristic scoring, including the new Greedy-Intersections strategy. The method runs many independent common subexpression elimination processes in parallel, exploring the search space through random pair substitutions and diverse selection strategies while sharing promising partial solutions. Tested on 164 ternary-coefficient schemes, the method achieves lower addition counts than the state-of-the-art Greedy-Potential on 103 schemes, matches it on 59, and is outperformed on 2. For most schemes, it gives equal or better results while being much faster, making it practical for algorithm exploration. All software and results are open source.
Keywords Fast matrix multiplication $\cdot$ Additive complexity reduction $\cdot$ Common subexpression elimination
@@ 1 Introduction
Recent work, such as the extensive study by Martensson et al (maartensson2025number), has established benchmarks and introduced methods like “Greedy-Potential” to automate this process. However, the CSE problem for matrix multiplication remains difficult. Greedy strategies often get stuck in local optima, while methods that exactly evaluate each substitution are slow for large schemes.
This paper presents a new approach using heuristic search and massive parallelism, focused on schemes with integer (especially ternary) coefficients. Two implementations are developed: a GPU-accelerated version that runs thousands of processes concurrently for maximum exploration, and a CPU version using OpenMP that runs tens to hundreds of processes for practical use. Both use the same core heuristic strategies. The main contributions are:
@@
  • A parallel optimization tool available in both GPU (CUDA) and CPU (OpenMP) versions. The tool runs many independent CSE processes concurrently. Processes use different strategies (rarely including Greedy-Potential as one of them) and share promising partial solutions, allowing broad exploration of the search space.
  • Comprehensive evaluation on 164 known ternary-coefficient ($Z_{T}=\{-1,0,1\}$) schemes. The method achieves lower addition counts than the fmm_add_reduction (fmm_add_reduction2025) baseline on 103 schemes, matches it on 59, and is outperformed on 2. All software and results are open source.
@@ 2.1 Additive Complexity in Fast Matrix Multiplication
For two schemes with the same rank, the one with fewer additions will execute faster. Therefore, after finding a good low-rank scheme, applying common subexpression elimination (CSE) to minimize its additive cost becomes a critical optimization step for practical implementation. The naive addition count serves as a starting point, while the optimized count, achieved through CSE, represents the real operation cost. For the irregular expression structures in fast matrix multiplication, CSE is a challenging combinatorial optimization problem.
@@ 2.2 Prior Work on Additive Minimization
The systematic study of additive minimization has a rich history, particularly for the foundational 2×2×2 case. Strassen’s original algorithm required 18 additions (strassen1969gaussian). Winograd reduced this to 15 (winograd1971multiplication), which was long believed optimal until Karstadt and Schwartz demonstrated that a change of basis could further reduce the count to 12 (karstadt2020matrix). This breakthrough showed that the framework for optimality itself could be expanded, yielding significant practical speedups. The approach of finding new, sparser algebraic representations via basis change, as further developed by Beniamini et al. (beniamini2020sparsifying) and Holtz et al. (holtz2025alternative), constitutes one major strategy for additive reduction.
The second major strategy is common subexpression elimination (CSE), which operates directly on a fixed scheme’s computational graph. The state-of-the-art in automated CSE is the Greedy-Potential algorithm introduced by Martensson et al. (maartensson2025number). This method, implemented in the fmm_add_reduction tool, scores a candidate substitution by its ”potential”: the number of new substitution possibilities created after applying it. This calculation requires a full trial substitution and re-analysis of all two-element subexpressions, which becomes computationally expensive for large schemes.
These strategies represent distinct approaches to the additive complexity problem. Recent work by Stapleton (stapleton202560) demonstrates a different direction: instead of optimizing additions for an existing scheme, a neural network with ternary regularization discovers a new low-rank tensor decomposition from scratch. The resulting scheme is then post-processed with Greedy-Potential CSE, achieving a record of 60 additions for a rank-23 scheme. A similar principle—searching for better algebraic forms before CSE optimization—underlies the flip-graph (kauers2023flip) exploration in previous work, where ternary flips (perminov2025fast) were used to find schemes with minimal naive addition counts before applying CSE.
This paper focuses on improving the CSE strategy itself. While Greedy-Potential is effective, its need for expensive calculations restricts the search scope. The proposed approach instead uses parallel heuristic search. It replaces exact scoring with faster approximate heuristics, including a new Greedy-Intersections strategy. This avoids computational bottlenecks, enabling broader exploration of the search space.
@@ 3.1 Matrix Multiplication Schemes
and reconstructs the transposed result matrix $C$, where $C=AB$, as:
@@
The tensors $(U,V,W)$ must satisfy the Brent equations (brent1970algorithms). A scheme is denoted by its format and rank as $(m,n,p:r)$. This work focuses on schemes where all coefficients in $U,V,W$ are restricted to the set $\{-1,0,1\}$ (ternary coefficients, denoted $Z_{T}$), which are particularly efficient for hardware implementation.
@@ 3.5.3 Dynamic Scheme Transformation via Random Flips
As an additional exploration mode, the GPU tool can incorporate transformations from the ternary flip graph (perminov2025fast). This mode was used during initial scheme discovery to find algebraic forms with potential for better additive reduction, though not in the final comparative experiments. The mode operates as follows:
@@ 4.1 Experimental Setup
The proposed parallel optimization tool is evaluated on 164 fast matrix multiplication schemes with ternary coefficients ($Z_{T}=\{-1,0,1\}$). The schemes include both previously known algorithms and new discoveries obtained through ternary flip graph exploration (perminov2025fast). Starting schemes were selected to have minimal naive addition counts as favorable starting points for additive reduction.
The baseline for comparison is the fmm_add_reduction tool (fmm_add_reduction2025). Three configurations of this baseline are tested:
  • Greedy-Vanilla: simple greedy strategy;
  • Greedy-Potential (5 steps): default greedy-potential mode with 5 values of the $\alpha$ parameter (linearly spaced from 0 to 0.5);
  • Greedy-Potential (40 steps): greedy-potential with 40 values of $\alpha$ (linearly spaced from 0 to 0.5), representing a more exhaustive search.
@@ 4.2 Additive Complexity Reduction Results
The proposed parallel optimization tool is compared against three baselines: Greedy-Vanilla (GV), Greedy-Potential with 5 $\alpha$ steps (GP-5), and Greedy-Potential with 40 $\alpha$ steps (GP-40). Across 164 ternary-coefficient schemes, the results fall into three categories.
4.2.1 Improved Addition Counts
Table 1 shows schemes where the proposed method achieves lower addition counts than GP-40. These 103 schemes demonstrate the effectiveness of the parallel heuristic search in escaping local optima.
Table 1: Schemes with improved addition counts (proposed method better than GP-40)
@@ 4.2.1 Improved Addition Counts
$m\times n\times p$ 5 40 $m\times n\times p$ 5 40
$2\times 3\times 10$ 50 198 135 135 135 134 $4\times 4\times 11$ 134 1150 534 528 524 519
$2\times 3\times 13$ 65 256 170 170 170 169 $4\times 4\times 12$ 145 1495 619 606 596 590
$2\times 4\times 7$ 45 308 181 178 177 174 $4\times 4\times 13$ 157 1474 650 638 635 628
$2\times 4\times 8$ 51 354 197 191 190 188 $4\times 4\times 14$ 169 1596 711 702 694 691
$2\times 4\times 9$ 59 309 210 208 208 207 $4\times 4\times 15$ 181 1679 741 738 735 729
$2\times 4\times 11$ 71 430 275 271 268 265 $4\times 4\times 16$ 192 2059 730 728 724 718
$2\times 4\times 12$ 77 484 282 277 276 274 $4\times 5\times 6$ 90 1023 401 393 386 380
$2\times 4\times 13$ 84 595 323 316 316 313 $4\times 5\times 7$ 104 931 423 413 405 402
$2\times 4\times 14$ 90 616 326 320 318 313 $4\times 5\times 8$ 118 1521 543 538 522 517
$2\times 4\times 15$ 96 662 366 358 357 351 $4\times 5\times 9$ 137 1217 564 555 547 542
$2\times 4\times 16$ 102 708 352 344 342 338 $4\times 5\times 10$ 151 1207 590 580 571 566
$2\times 5\times 5$ 40 283 156 154 154 153 $4\times 5\times 11$ 165 1801 718 702 689 681
$2\times 5\times 6$ 47 332 189 186 184 181 $4\times 5\times 12$ 179 1977 778 772 761 750
$2\times 5\times 7$ 57 340 197 194 192 189 $4\times 6\times 6$ 105 894 467 435 435 430
$2\times 5\times 8$ 65 376 222 218 218 214 $4\times 6\times 7$ 123 1586 535 533 518 517
$2\times 5\times 9$ 72 465 275 270 270 266 $4\times 6\times 8$ 140 1248 576 558 558 551
$2\times 5\times 11$ 87 540 331 326 326 323 $4\times 6\times 9$ 160 1472 720 651 647 646
$2\times 5\times 12$ 94 664 336 330 328 322 $4\times 6\times 10$ 175 1878 759 738 723 715
$2\times 6\times 7$ 68 396 233 230 230 226 $4\times 7\times 7$ 145 1381 680 646 632 629
$2\times 6\times 8$ 77 456 272 269 269 266 $4\times 7\times 8$ 164 1505 752 729 698 690
$2\times 6\times 9$ 86 548 302 296 295 293 $4\times 7\times 9$ 187 2059 806 800 788 785
$2\times 6\times 10$ 94 668 334 327 327 325 $4\times 8\times 8$ 182 1884 861 812 800 795
$2\times 7\times 8$ 90 648 350 333 331 329 $5\times 5\times 5$ 93 843 392 391 387 384
$2\times 7\times 9$ 102 678 398 384 382 379 $5\times 5\times 6$ 110 1215 491 479 466 460
$3\times 3\times 6$ 43 284 168 164 164 160 $5\times 5\times 7$ 127 1607 561 553 539 531
$3\times 3\times 9$ 65 347 218 216 216 215 $5\times 5\times 8$ 144 1924 645 638 631 620
$3\times 3\times 12$ 86 582 303 297 297 289 $5\times 5\times 9$ 167 1814 706 699 690 682
$3\times 3\times 14$ 101 664 373 365 363 361 $5\times 5\times 10$ 184 2116 796 787 781 774
$3\times 4\times 6$ 56 359 211 210 210 209 $5\times 5\times 11$ 202 2272 879 872 859 850
$3\times 4\times 10$ 93 582 351 346 346 345 $5\times 5\times 12$ 220 2444 848 835 811 799
$3\times 4\times 15$ 138 958 544 537 535 533 $5\times 6\times 6$ 130 1716 589 587 573 562
$3\times 5\times 5$ 58 357 230 224 223 221 $5\times 6\times 7$ 150 2039 696 691 671 664
$3\times 5\times 6$ 70 561 270 267 267 265 $5\times 6\times 8$ 170 2312 747 740 728 720
$3\times 5\times 7$ 83 494 315 308 305 303 $5\times 6\times 9$ 197 2376 884 875 850 840
$3\times 5\times 9$ 105 646 390 384 384 383 $5\times 6\times 10$ 217 2772 956 954 940 925
$3\times 5\times 10$ 116 714 402 391 391 389 $5\times 7\times 7$ 176 2605 832 820 799 788
$3\times 5\times 11$ 128 967 497 489 488 484 $5\times 7\times 8$ 206 1880 948 892 882 879
$3\times 5\times 12$ 140 1152 491 483 483 482 $5\times 7\times 9$ 231 2554 993 981 971 963
$3\times 6\times 6$ 85 998 367 364 359 352 $5\times 8\times 8$ 230 2747 1019 1010 980 974
$3\times 6\times 7$ 100 704 364 352 351 350 $6\times 6\times 6$ 153 2182 704 685 671 655
$3\times 6\times 8$ 113 1240 467 459 446 445 $6\times 6\times 7$ 183 2502 810 790 777 769
$3\times 6\times 9$ 127 1095 519 502 493 489 $6\times 6\times 8$ 203 1994 896 880 868 836
$3\times 6\times 10$ 140 1152 494 484 477 476 $6\times 6\times 9$ 225 2440 1029 951 936 923
$3\times 7\times 7$ 115 789 441 434 434 432 $6\times 7\times 7$ 215 2004 965 900 886 880
$3\times 7\times 9$ 147 1002 592 583 580 579 $6\times 7\times 8$ 239 2263 1112 1050 1042 1027
$4\times 4\times 4$ 49 474 167 164 163 159 $6\times 7\times 9$ 268 3062 1184 1166 1152 1146
$4\times 4\times 5$ 61 452 247 237 237 235 $6\times 8\times 8$ 266 2780 1244 1214 1181 1161
$4\times 4\times 6$ 73 540 288 282 281 280 $7\times 7\times 7$ 250 2417 1119 1077 1070 1067
$4\times 4\times 7$ 85 631 326 323 320 319 $7\times 7\times 8$ 279 2926 1369 1260 1248 1237
$4\times 4\times 8$ 96 973 396 391 387 377 $7\times 7\times 9$ 316 3452 1460 1404 1385 1383
$4\times 4\times 9$ 110 925 445 424 422 419 $8\times 8\times 8$ 343 4434 1748 1709 1668 1661
$4\times 4\times 10$ 122 879 493 482 482 480

### v2

@@ Abstract
This paper presents a parallel random-search method for reducing additive complexity in fast matrix multiplication algorithms with ternary coefficients $\{-1,0,1\}$. The approach replaces expensive exact evaluation with fast heuristic scoring, including the new Greedy-Intersections strategy. The method runs many independent common subexpression elimination processes in parallel, exploring the search space through random pair substitutions and diverse selection strategies while sharing promising partial solutions. Tested on 149 ternary-coefficient schemes, the method achieves lower addition counts than the state-of-the-art Greedy-Potential on 102 schemes (including 57 new best-known results for optimal-rank schemes), matches it on 45, and is outperformed on only 2. For most schemes, it provides equal or better results while being significantly faster, making it practical for algorithm exploration. All software and results are open source.
Keywords Fast matrix multiplication $\cdot$ Additive complexity reduction $\cdot$ Common subexpression elimination $\cdot$ Ternary coefficients set
@@ 1 Introduction
This work focuses exclusively on schemes with coefficients in the ternary set $Z_{T}=\{-1,0,1\}$. Ternary schemes are particularly valuable for hardware implementation due to their simple coefficient representation, and they are also well-suited for testing addition reduction methods. The restriction to $Z_{T}$ allows for efficient data representations and specialized heuristics while maintaining relevance to practical fast matrix multiplication.
Recent work, such as the extensive study by Martensson et al [maartensson2025number], has established benchmarks and introduced methods like “Greedy-Potential” to automate this process. However, the CSE problem for matrix multiplication remains difficult. Greedy strategies often get stuck in local optima, while methods that exactly evaluate each substitution are slow for large schemes.
This paper presents a parallel heuristic approach to additive complexity reduction for ternary-coefficient schemes. The method addresses the combinatorial challenge by running many independent CSE processes concurrently, each guided by different selection strategies. Two implementations are developed: a GPU-accelerated version that runs thousands of processes concurrently for maximum exploration, and a CPU version using OpenMP that runs tens to hundreds of processes for practical use. Both use the same core heuristic strategies.
The main contributions are:
@@
  • A parallel optimization tool available in both GPU (CUDA) and CPU (OpenMP) versions. The tool runs many independent CSE processes concurrently. Processes use different strategies and share promising partial solutions, allowing wide exploration of the search space.
  • Comprehensive evaluation on 149 known ternary-coefficient schemes. The method achieves lower addition counts than the fmm_add_reduction [fmm_add_reduction2025] baseline on 102 schemes, matches it on 45, and is outperformed on 2.
  • For 57 schemes with optimal known rank, the method achieves new reference addition counts, advancing the state-of-the-art for practical matrix multiplication implementations. All software and results are open source.
@@ 2.1 Additive Complexity in Fast Matrix Multiplication
For two schemes with the same rank, the one with fewer additions will generally execute faster, provided other factors like memory access patterns and cache utilization are comparable. In practical implementations, these architectural considerations also affect runtime, but they are beyond the scope of this paper, which focuses exclusively on minimizing the arithmetic operation count. Therefore, after finding a good low-rank scheme, applying common subexpression elimination (CSE) to minimize its additive cost becomes a critical optimization step. The naive addition count serves as a starting point, while the optimized count, achieved through CSE, represents the real arithmetic cost. For the irregular expression structures in fast matrix multiplication, CSE is a challenging combinatorial optimization problem where even small improvements are valuable, especially for schemes with an already optimal rank, as they yield new state-of-the-art algorithms.
@@ 2.2 Prior Work on Additive Minimization
The systematic study of additive minimization has a rich history, particularly for the foundational $2\times 2\times 2$ case. Strassen’s original algorithm required 18 additions [strassen1969gaussian]. Winograd reduced this to 15 [winograd1971multiplication], which was long believed optimal until Karstadt and Schwartz demonstrated that a change of basis could further reduce the count to 12 [karstadt2020matrix]. This breakthrough showed that the framework for optimality itself could be expanded, yielding significant practical speedups. The approach of finding new, sparser algebraic representations via basis change, as further developed by Beniamini et al. [beniamini2020sparsifying] and Holtz et al. [holtz2025alternative], constitutes one major strategy for additive reduction.
The second major strategy is common subexpression elimination (CSE), which operates directly on a fixed scheme’s computational graph. Tools like PLinOpt [dumas2024place] provide a general framework for optimizing linear, bilinear, and trilinear programs, including the generation of compact straight-line programs from matrix multiplication schemes. This work handles general integer and rational coefficients and can produce in-place accumulation programs. For the specific and highly structured problem of minimizing additions in ternary-coefficient schemes, more specialized tools have been developed. The state-of-the-art in automated CSE for this domain is the Greedy-Potential algorithm introduced by Martensson et al. [maartensson2025number]. This method, implemented in the fmm_add_reduction tool [fmm_add_reduction2025], scores a candidate substitution by its ”potential”: the number of new substitution possibilities created after applying it. This calculation requires a full trial substitution and re-analysis of all two-element subexpressions, which becomes computationally expensive for large schemes.
These strategies represent distinct approaches to the additive complexity problem. Recent work by Stapleton [stapleton202560] demonstrates a different direction: instead of optimizing additions for an existing scheme, a neural network with ternary regularization discovers a new low-rank $3\times 3\times 3$ tensor decomposition from scratch. The resulting scheme is then post-processed with Greedy-Potential CSE, achieving a record of 60 additions for a rank-23 scheme. A similar principle – searching for better algebraic forms before CSE optimization – underlies the flip-graph [kauers2023flip] exploration in previous work, where ternary flips [perminov2025fast] were used to find schemes with minimal naive addition counts before applying CSE.
This paper focuses on improving the CSE strategy itself. While Greedy-Potential is effective, its need for expensive calculations restricts the search scope. The proposed approach instead uses parallel heuristic search. It replaces exact scoring with faster approximate heuristics, including a new Greedy-Intersections strategy. This avoids computational bottlenecks, enabling wider exploration of the search space.
@@ 3.1 Matrix Multiplication Schemes
and reconstructs the result matrix $C$, where $C=AB$, as:
@@
The tensors $(U,V,W)$ must satisfy the Brent equations [brent1970algorithms]. A scheme is denoted by its format and rank as $(m,n,p:r)$. This work focuses on schemes where all coefficients in $U,V,W$ are restricted to the set $\{-1,0,1\}$, which are particularly efficient for hardware implementation.
@@ 3.5.3 Dynamic Scheme Transformation via Random Flips
As an additional exploration mode, the GPU tool can incorporate transformations from the ternary flip graph [perminov2025fast]. This mode was used during initial scheme discovery to find algebraic forms with potential for better additive reduction, though not in the final comparative experiments. The mode operates as follows:
@@ 4.1 Experimental Setup
The proposed parallel optimization tool is evaluated on 149 fast matrix multiplication schemes with ternary coefficients. The schemes include both previously known algorithms and new discoveries obtained through ternary flip graph exploration [perminov2025fast]. Starting schemes were selected to have minimal naive addition counts as favorable starting points for additive reduction.
The baseline for comparison is the fmm_add_reduction tool [fmm_add_reduction2025]. Three configurations of this baseline are tested:
  • Greedy-Vanilla: simple greedy strategy (GV);
  • Greedy-Potential (5 steps): default greedy-potential mode with 5 values of the $\alpha$ parameter linearly spaced from 0 to 0.5 (GP-5);
  • Greedy-Potential (40 steps): greedy-potential with 40 values of $\alpha$ linearly spaced from 0 to 0.5, representing a more exhaustive search (GP-40).
@@ 4.2 Additive Complexity Reduction Results
Across 149 ternary-coefficient schemes, the results demonstrate two key outcomes: the establishment of new state-of-the-art addition counts for optimal-rank schemes, and a general ability to find better reductions than exhaustive search methods across a wide set of problems.
4.2.1 New State-of-the-Art Addition Counts
A primary practical achievement is the improvement of 57 matrix multiplication schemes that already have the best-known (optimal) rank for their dimensions. For these algorithms, which represent a multiplicative complexity bound, proposed method finds a more efficient evaluation scheme, setting a new reference for their total arithmetic cost. Table 1 lists of these new SOTA results.
Table 1: Schemes with optimal rank achieving new SOTA addition counts
@@ 4.2.1 New State-of-the-Art Addition Counts
$m\times n\times p$ $r$ 5 40 $m\times n\times p$ $r$ 5 40
$2\times 3\times 10$ 50 198 135 135 135 134 $4\times 6\times 7$ 123 1586 535 533 518 517
$2\times 3\times 13$ 65 256 170 170 170 169 $4\times 6\times 8$ 140 1248 576 558 558 551
$2\times 4\times 7$ 45 308 181 178 177 174 $4\times 6\times 10$ 175 1878 759 738 723 715
$2\times 4\times 8$ 51 354 197 191 190 188 $4\times 7\times 8$ 164 1505 752 729 698 690
$2\times 4\times 9$ 59 309 210 208 208 207 $4\times 8\times 8$ 182 1884 861 812 800 795
$2\times 4\times 11$ 71 430 275 271 268 265 $5\times 5\times 5$ 93 843 392 391 387 383
$2\times 4\times 12$ 77 484 282 277 276 274 $5\times 5\times 6$ 110 1215 491 479 466 460
$2\times 4\times 14$ 90 616 326 320 318 313 $5\times 5\times 7$ 127 1607 561 553 539 531
$2\times 4\times 15$ 96 662 366 358 357 351 $5\times 5\times 8$ 144 1924 645 638 631 620
$2\times 4\times 16$ 102 708 352 344 342 338 $5\times 5\times 9$ 167 1814 706 699 690 682
$2\times 5\times 5$ 40 283 156 154 154 153 $5\times 5\times 10$ 184 2116 796 787 781 774
$2\times 5\times 6$ 47 332 189 186 184 181 $5\times 5\times 11$ 202 2272 879 872 859 850
$2\times 5\times 9$ 72 465 275 270 270 266 $5\times 5\times 12$ 220 2444 848 835 811 799
$2\times 5\times 11$ 87 540 331 326 326 323 $5\times 6\times 6$ 130 1716 589 587 573 562
$2\times 5\times 12$ 94 664 336 330 328 322 $5\times 6\times 7$ 150 2039 696 691 671 664
$2\times 6\times 9$ 86 548 302 296 295 293 $5\times 6\times 8$ 170 2312 747 740 728 720
$2\times 6\times 10$ 94 668 334 327 327 325 $5\times 6\times 9$ 197 2376 884 875 850 840
$3\times 5\times 5$ 58 357 230 224 223 221 $5\times 6\times 10$ 217 2772 956 954 940 925
$4\times 4\times 5$ 61 452 247 237 237 233 $5\times 7\times 7$ 176 2605 832 820 799 788
$4\times 4\times 6$ 73 540 288 282 281 280 $5\times 8\times 8$ 230 2747 1019 1010 980 974
$4\times 4\times 7$ 85 631 326 323 320 319 $6\times 6\times 6$ 153 2182 704 685 671 655
$4\times 4\times 8$ 96 973 396 391 387 377 $6\times 6\times 7$ 183 2502 810 790 777 769
$4\times 5\times 6$ 90 1023 401 393 386 380 $6\times 6\times 8$ 203 1994 896 880 868 836
$4\times 5\times 7$ 104 931 423 413 405 400 $6\times 6\times 9$ 225 2440 1029 951 936 923
$4\times 5\times 8$ 118 1521 543 538 522 513 $6\times 7\times 7$ 215 2004 965 900 886 880
$4\times 5\times 10$ 151 1207 590 580 571 566 $6\times 7\times 8$ 239 2263 1112 1050 1042 1027
$4\times 5\times 11$ 165 1801 718 702 689 681 $6\times 7\times 9$ 268 3062 1184 1166 1152 1146
$4\times 5\times 12$ 179 1977 778 772 761 750 $6\times 8\times 8$ 266 2780 1244 1214 1181 1161
$4\times 6\times 6$ 105 894 467 435 435 430
4.2.2 General Improvements Across All Schemes
Beyond establishing new SOTA records, the parallel heuristic search demonstrates high effectiveness in the additive minimization task. It finds a lower addition count than the exhaustive GP-40 method for an additional 45 schemes that do not have optimal rank. Table 2 shows a representative sample of these improvements, confirming the method’s ability to escape local optima encountered by other strategies.
Table 2: Representative general improvements (non-optimal rank schemes)
Format Rank Naive GV GP GP Prop. Format Rank Naive GV GP GP Prop.
$m\times n\times p$ $r$ 5 40 $m\times n\times p$ $r$ 5 40
$2\times 4\times 13$ 84 595 323 316 316 313 $3\times 6\times 10$ 140 1152 494 484 477 476
$2\times 5\times 7$ 57 340 197 194 192 189 $3\times 7\times 7$ 115 789 441 434 434 432
$2\times 5\times 8$ 65 376 222 218 218 214 $3\times 7\times 9$ 147 1002 592 583 580 579
$2\times 6\times 7$ 68 396 233 230 230 226 $4\times 4\times 4$ 49 474 167 164 163 159
$2\times 6\times 8$ 77 456 272 269 269 266 $4\times 4\times 9$ 110 925 445 424 422 419
$2\times 7\times 8$ 90 648 350 333 331 329 $4\times 4\times 10$ 122 879 493 482 482 480
$2\times 7\times 9$ 102 678 398 384 382 379 $4\times 4\times 11$ 134 1150 534 528 524 519
$3\times 3\times 6$ 43 284 168 164 164 160 $4\times 4\times 12$ 145 1495 619 606 596 590
$3\times 3\times 9$ 65 347 218 216 216 215 $4\times 4\times 13$ 157 1474 650 638 635 628
$3\times 3\times 12$ 86 582 303 297 297 289 $4\times 4\times 14$ 169 1596 711 702 694 691
$3\times 3\times 14$ 101 664 373 365 363 361 $4\times 4\times 15$ 181 1679 741 738 735 729
$3\times 4\times 6$ 56 359 211 210 210 209 $4\times 4\times 16$ 192 2059 730 728 724 718
$3\times 4\times 10$ 93 582 351 346 346 345 $4\times 5\times 9$ 137 1217 564 555 547 542
$3\times 4\times 15$ 138 958 544 537 535 533 $4\times 6\times 9$ 160 1472 720 651 647 646
$3\times 5\times 6$ 70 561 270 267 267 265 $4\times 7\times 7$ 145 1381 680 646 632 629
$3\times 5\times 7$ 83 494 315 308 305 303 $4\times 7\times 9$ 187 2059 806 800 788 785
$3\times 5\times 9$ 105 646 390 384 384 383 $5\times 7\times 8$ 206 1880 948 892 882 879
$3\times 5\times 11$ 128 967 497 489 488 484 $5\times 7\times 9$ 231 2554 993 981 971 963
$3\times 5\times 12$ 140 1152 491 483 483 482 $7\times 7\times 7$ 250 2417 1119 1077 1070 1067
$3\times 6\times 6$ 85 998 367 364 359 352 $7\times 7\times 8$ 279 2926 1369 1260 1248 1237
$3\times 6\times 7$ 100 704 364 352 351 350 $7\times 7\times 9$ 316 3452 1460 1404 1385 1383
$3\times 6\times 8$ 113 1240 467 459 446 445 $8\times 8\times 8$ 343 4434 1748 1709 1668 1661
$3\times 6\times 9$ 127 1095 519 502 493 489
In total, the proposed method achieves a strict improvement over GP-40 for 102 of the 149 tested schemes (68.4%). The complete list of all 102 improved schemes is available in the included open-source repository.